16. Calculate the number of moles for 11.2 liters of fluorine, F2 (in K.N.) and for 35.5 g of chlorine, Cl2 (Ar = 35.5)


1
Expert's answer
2021-06-04T16:19:05-0400

Solution:

At STP, one mole of any gas occupies a volume of 22.4 L.

Thus, (11.2 L F2) × (1 mol / 22.4 L) = 0.5 mol F2

Moles of fluorine gas (F2) = 0.5 mol


The molar mass of Cl2 is 2 × Ar(Cl) = 2 × 35.5 = 71 g/mol.

Thus, (35.5 g Cl2) × (1 mol Cl2 / 71 g Cl2) = 0.5 mol Cl2

Moles of chlorine gas (Cl2) = 0.5 mol


Answer:

Moles of fluorine gas (F2) = 0.5 mol

Moles of chlorine gas (Cl2) = 0.5 mol